# From the CDC to emerging infectious disease publics: The long-now of polarizing and complex health crises

### v2

@@ Abstract.
As the COVID-19 pandemic evolved, the Center for Disease Control and Prevention used Twitter to share updates about the virus and safety guidelines, reaching millions instantly, in what we call the CDC public. We analyze two years of tweets, from, to, and about the CDC using a mixed-methods approach to characterize the nature and credibility of COVID-19 discourse and audience engagement. We found that the CDC is not engaging in two-way communication with the CDC publics and that discussions about COVID-19 reflected societal divisions and political polarization. We introduce a crisis message journey concept showing how the CDC public responds to the changing nature of the crisis (e.g., new variants) using “receipts” of earlier, and at times contradictory, guidelines. We propose design recommendations to support the CDC in tailoring messages to specific users and publics (e.g., users interested in racial equity) and in managing misinformation, especially in reaction to crisis flashpoints.
Twitter, Digital publics, Public health crisis, Crisis communication, Public policy discourse, Temporal crisis distance, affordances, change-point detection
copyright: acmlicensedjournal: TOCHIjournalyear: 2025journalvolume: XXjournalnumber: XXarticle: TOCHIpublicationmonth: 4doi: XX.XXXX/XXXXXXXccs: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
Twitter is a virtual “public square” (Jenzen et al., 2021) where different ad hoc “publics” can be identified using hashtags (Bruns and Moe, 2014) which are consumed by interested audiences (Livingstone, 2005). They can also be formed around mentions that are searchable by other users (Kountouri and Kollias, 2023). Retweets, quote-tweets, and responses to Tweets are associated with propagating messages in different ways. Given how different affordances of the platform (Faraj and Azad, [n. d.]; Leonardi and Vaast, [n. d.]) can affect the way frames are propagated through retweets and directed through mentions at other members of the public, we ask:
RQ1a: What are the properties of users whose messages are propagated the most?
Given that earlier work shows how the use of low-credibility information sources (Wang et al., 2023; Hobbs et al., 2024), the use of rich media in the message (Jin et al., 2016; Gupta et al., 2013), and the sentiment portrayed in a message (Valdez et al., 2020) all affect the way messages are propagated in a network, we ask:
RQ1b: How do the properties of a message affect the way they are propagated in the CDC public?
Finally, given that earlier work shows how the properties of the users disseminating the messages, specifically their follower-following ratio (signifying their influence in the network) and their verified status (Harris et al., 2024) affect the way their messages are propagated in publics, we ask:
RQ1c: How are messages from verified users and users with high follower-following ratios in the CDC public different from other users?
@@
RQ2a: What are the major frames of discussion in different publics and how are they polarized?
@@
RQ2b: Who are the main stakeholders in our public? How are they polarized around different issues?
Public health crises are often a time of uncertainty, confusion, and frustration. These heightened emotions amplify community discourse, particularly on social media platforms (Pine et al., 2021). This discourse can be split into two categories: communication pertaining to the risk of infection, and the risks associated with the authorities’ crisis management measures (Gui et al., 2018). As public health crises progress over time, the responses to the public health emergency change, both on the side of public health organizations like the CDC, and from publics (Coombs, 2021). This is especially true in specific flashpoints like when the first health worker is infected by a patient. Given the importance of changing messaging as the temporal crisis distance increases (Kim and Jin, 2020; Kim, 2022), especially around flashpoints (Dalrymple et al., 2016), we ask:
RQ3a: What are the COVID flashpoints in the CDC public?
Some users will have more of an effect on the discourse on CDC publics over time. To better understand what makes for a more influential messenger over time on the CDC public, we ask:
RQ3b: What are the properties of users most committed to the CDC public?
@@
We adapt the public policy discourse framing analytical framework described in Park and Lee (Park and Lee, 2020)[P.641] to model how Twitter is used as a communication channel to and from the CDC throughout a two-year public health emergency - COVID (see Figure 1). The framework suggests that public policy comprises communication between the state and society (public & private). In this framework, “communicative interaction among political actors operates in a multilayered policy landscape…[to advocate] for ”policy goals (in our case, COVID mitigation policy) and formulating the policy tools (e.g., mask and vaccine mandates) needed to achieve these goals. This communication is mediated through either a one-way or two-way channel to the public. On the other hand, the public will adopt policies and support policy tools depending on their needs and persuasions as stakeholders.
Our study shows that the affordances of technology platforms used in communication need to be taken into account when creating a public for public health crises. We found that indeed, most of the frames in the CDC public were affectively (positive vs. negative) and ideologically (liberal vs. conservative) polarized. The most propagated frames focused on polarizing topics including vaccinations, reopening schools, and arguing about whether the pandemic is a hoax. The mentions network was sparse with strong mutual connections, meaning that echo chambers were formed where some members corresponded with each other while not mentioning the others. Users sharing low-credibility sources were less likely to receive mentions, they were more likely to mention others. This shows that they were more committed to sharing their message in the network. The more low-credibility sources are used by an account, the less likely they are to receive mentions from other accounts. However, we also found that verified accounts, which are more likely to be retweeted, also share low-credibility sources. We suggest new elements to the public policy discourse framing analytical framework based on Twitter affordances and user behavior. We also provide design recommendations for future public policy communication that incorporate emerging technologies, especially generative AI. Finally, we introduce the concept of a crisis messaging journey and show how public health organizations like the CDC need to control messaging about emerging infectious diseases for long time horizons. In Figure 1, we show two main flashpoints, one a sentiment change-point, the other focused on how new variants forced in a change in messaging (more are discussed in §7).
@@ 2. Related Work
In this section, we start by reviewing Twitter affordances, namely managing audiences and association with content. Next, we analyze digital publics, focusing on polarizing issues. Finally, we review digital communication in public health crises.
2.1. Twitter affordances and audience management
Technological affordances are defined as the relationship between users and ”their perceptions of environments” (Pavelko and Myrick, [n. d.], P.361). Affordances can explain the influence of different technologies on human behavior (Faraj and Azad, [n. d.]; Leonardi and Vaast, [n. d.]). Twitter has numerous platform-specific affordances that enable people to communicate with each other, associate with content, and provide or receive direct (DeVito et al., [n. d.]) or indirect feedback (Randazzo and Ammari, 2023) due to the searchability and permanence of the content (Leonardi and Vaast, [n. d.]). We will focus on two main affordances: (1) hashtags and (2) interactions including mentions, retweets (Kim and Yoo, 2012; Boyd et al., 2010), likes (McDonald, 2021; Huang et al., 2018; Wang et al., 2016), replies (Kim and Yoo, 2012), and quote-tweets (Garimella et al., 2016; Zade et al., 2024).
2.1.1. Imagined audiences on Twitter
Hashtags are keywords that allow users to search for or find a topic of interest. This allows users to build an imagined audience (Litt, 2012) whom they believe to be interested in a particular topic, “to mark participation, assert individual identity, promote group identity, and support or challenge a frame.” (Stewart et al., 2017) For example, hashtags make issues of interest like Black Lives Matter (#BLM) and #stopasianhate more searchable and accessible, which in turn allows users to reach wider audiences (Tong et al., 2022). Additionally, Huang et al. (Huang et al., 2010) argue that hashtags are conversational tools that motivate people to share their thoughts on the topic of interest.
Another way to build an audience is by using mentions (e.g., @user). When a Tweet mentions a user, only the target user receives a notification, but other users can view and search for the Tweet. In other words, users searching for Tweets made by an account will also see tweets mentioning that account. This means one can expand their audience by targeting an account belonging to a person (e.g., politician or journalist) or organization central to the discourse on a particular topic (McGregor, 2019). Users will typically write negative Tweets mentioning a specific account to send a hostile message rather than start a discussion with the target user (Hemsley et al., 2018) with the hope that a provocative message to an influential node in the network would increase the size of their audience.
Retweets propagate messages throughout the Twitter network, and whether intended or not, serve as endorsements of the original Tweet message (Kim and Yoo, 2012) and are thus associated with the virality of a message (Bonsón et al., 2019). Earlier work trained models using the number of retweets to better understand more influential messages in the network (c.f.,(Jenders et al., 2013; Suh et al., 2010; Bakshy et al., 2011)). Likes also propagate messages to a more targeted audience than retweets (Rathnayake and Suthers, 2017; Sekimoto et al., 2020). Thus, they show the popularity of the message since the user “liking” the Tweet is an implicit approval of the message (Bonsón et al., 2019; Sekimoto et al., 2020).
Responses to Tweets are usually associated with commitment to an issue (Bonsón et al., 2019) as it provides a chance to re-frame the conversation (Zade et al., 2024) by providing a positive or negative view, thus pushing the user’s view of the original message (Garimella et al., 2016). Quote-tweets, much like forwarded messages, not only allow the user to push the original message in a new direction, but also “broadcasts” (Zade et al., 2024) this new contextualized view to networks in which they are embedded (Garimella et al., 2016).
2.1.2. Rich Media on Twitter
While social media was traditionally more text-heavy, social media sites like Twitter, Instagram, and TikTok afford users the capacity to create more rich content - including images and videos. This becomes especially useful on Twitter given the words limits of the platform. Adding rich media allows users to extend their text posts (Jin et al., 2016; Gupta et al., 2013).
Shi et al. show that rich media increases the likelihood that users will propagate messages (e.g., through likes and retweets) (Shi et al., 2018; Jiang et al., 2019) and increase the perceived credibility of the message (Yin and Zhang, 2020; Li et al., 2022). Richer messages also increase the persuasiveness (Zhou et al., 2021b) and the reach (Wang et al., 2023) of COVID-19 misinformation.
2.1.3. Verified Twitter accounts
Verified Twitter accounts are central to the mechanism by which users search for and propagate information on the platform (González-Bailón and De Domenico, 2021; Paul et al., 2019) especially in the event of a crisis like COVID (Zhou et al., 2023). In fact, Twitter shared that it would increase the number of local verified accounts sharing information about COVID (Support, 2020) to provide more trusted information in a timely manner (Lunden, 2020). One example Zhou et al. (Zhou et al., 2023, P.22) provide is Dr. Eric Ding (@DrEricDing) who became a verified account on January 25, 2020. Additionally, self-disclosure by verified accounts about their COVID challenges increased the change of self-disclosure in the network (Du et al., 2022).
However, verified accounts can also become “perceived experts” who spread misinformation about vaccines (Harris et al., 2024) especially in regard to the side-effects of new COVID vaccines like the Oxford/AstraZeneca (Hobbs et al., 2024).
@@ 2.2. Digital Publics and Crisis Management
In these public spheres, stakeholders come together to talk about issues (c.f., (Dewey, 1927; Rauchfleisch et al., 2021)) defined as “socially constructed matters of public interest around which there is contention” (Stoltenberg, 2024, P.2009). These publics evolve over time around specific topics of interest signified by hashtags that convey the content of the messages as described in §2.1 (Bruns and Burgess, 2011).
@@ 2.2.1. Hashtags, mentions, and publics
2.2.2. Framing in Crisis Publics
@@
Organizations like FEMA and the CDC aim to frame their messages in a way that counters politicization and instead, focuses on their missions (Vanessa Murphree and Blevens, 2009). Earlier work suggests that these organizations should consider a crisis from different facets, such as the perspectives of different publics, and find the appropriate framing for each (Stephens et al., 2005).
2.2.3. Misinformation and Distrust
Despite this information shared on social media channels being useful to many, increased communication from public health systems has also given rise to conspiracy theories and distrust. Kou et al. explain that conspiracy theories are a type of rumor. During the Zika virus outbreak, social media served as a source of up-to-date Zika-related news, as well as a hub for conspiracy theories about the virus (Kou et al., 2017). Kou et al. revealed that these theories were the public’s attempt to make sense of the significant shift in society caused by the Zika outbreak (Kou et al., 2017). This coping mechanism quickly becomes convoluted when health authorities do not have complete information about the virus causing the outbreak (Gui et al., 2017). The lack of trust in formal sources of crisis-related information spurs greater conspiracy theories. Those in vulnerable groups tend to experience higher distrust in public health and healthcare systems (Abelson et al., 2009). The HIV/AIDS conspiracy theories, for example, perpetuated the idea that the pandemic was planned by the government to target and harm specific populations (Kou et al., 2017). These theories often have significant consequences. In this example, distrust in the government led to AIDS patients refusing urgent testing and treatment (Casley, 2010). Similarly, images and memes were used to spread misinformation (Wang et al., 2023) and hate (Liao et al., 2023; Vishwamitra et al., 2024; Javed et al., 2023) in COVID mitigation discourse.
@@ 2.3. Digital Communication in Public Health Crises
Gui et al. propose including the public in the decision-making process behind crisis communication to reduce the distrust between crisis management organizations like the CDC and the public (Gui et al., 2018). An example of such engagement with the public is the organization of a Hackathon to brainstorm COVID-19 solutions in different social domains held in Germany (Haesler et al., 2021).
Usually however, public health agencies tend to utilize social media as a means of broadcasting crisis management information rather than a two-way communication tool (Graham, 2014). By altering this perspective and examining the public’s sentiment through social media interactions, information gaps and overlooked concerns can be addressed sooner, limiting distrust.
2.3.1. Emerging Infectious Diseases: Rolling uncertainty, risk, and public relations
Fox (Fox, 2012) suggests embracing the “inevitable uncertainty” of medicine where scientists and medical professionals have to contend with “certainty for now” instead of more objective certainty. This uncertainty is significantly higher for emerging infectious diseases (EIDs) like Ebola and COVID because the provision of medical expertise and the touting of existing protocols provides an “illusion of certainty.” (Dalrymple et al., 2016) For example, when the CDC protocols set to prevent the spread of Ebola failed, the CDC tended to emphasize what was already known about the disease (Dalrymple et al., 2016, P.455).
Earlier work shows that when organizations like the CDC emphasize certainty in a chaotic uncertain times, this amplifies the uncertainty that the general public feels, and increases doubts about the institutions as well (Dalrymple et al., 2016).
2.3.2. Crisis and Emergency Risk Communication
The Crisis and Emergency Risk Communication (CERC) model describes how, to better manage uncertainty at the height of crisis, an organization has to engage in two-way communication (Reynolds and Seeger, 2005) where the public can give feedback about the management of a crisis (Renn, 2015). Specifically, public facing organizations like the CDC or FEMA need to engage in what Dozier et al. (Waters and Jamal, 2011) call symetric two-way communication which provides for a mutual understanding between the public and organizations which could lead to changes in their communication strategies (Grunig and Grunig, 2008; Waters and Jamal, 2011). However, earlier work has shown that CDC’s online communication has been mostly top-down (from the CDC to users) instead of bottom-up communication (from users to the CDC; (Gesser-Edelsburg et al., 2014).
In the pre-crisis stage, the CERC model suggests that public health organizations should strive to “alliances and cooperation with agencies, organizations and groups.” (Reynolds and Seeger, 2005, P.52). After the initial crisis event (e.g., a novel virus infects a substantial part of the population), public health organizations are tasked with discussing the frame of disease detection (disease symptoms) and disease prevention (taking preventive measures) (Umphrey, 2003; Higgins et al., 2006). In the maintenance phase of the crisis, public health organizations need to communicate healthcare services (services provided by healthcare system), scientific advances like discovering new evidence about the disease, and lifestyle risk factors (e.g., personal habits associated with the disease) (Ngai et al., 2020). Assuming that there is a resolution to the crisis (e.g, effective vaccine is produced), the organization needs to evaluate the communication experience by engaging with the public (Reynolds and Seeger, 2005).
One major challenge faced by scholars lies in the lack of a mechanism to determine different phases of an outbreak (beginning, middle, and end) since they are not predictable, especially in IEDs (Dalrymple et al., 2016, P.457). We discuss the changes between these fluid phases in the next subsection.
2.3.3. Flashpoints of change in a health crisis
Events like the first transmission of a disease from a patient to a healthcare worker (e.g., Ebola spreading from patient to nurse), serve as “flashpoints of change” (Vanderford et al., 2007) that necessitate real-time “rapid redirection on social media.” (Dalrymple et al., 2016) Any such rapid change is challenged by social media affordances like persistence of earlier data shared by public health organizations (see §2.1).
Earlier work shows that as the temporal crisis distance increases, the sentiment of the crisis publics tends to be more positive, and their response more supporting of the crisis organization (e.g., FEMA or a drug company) (Kim and Jin, 2020; Kim, 2022). Though they did not focus on any crisis publics, Valdez et al. used topic modeling and sentiment analysis to measure public mood changes over a period of four months (starting in January of 2020 and ending in early April of 2020) (Valdez et al., 2020). The use of Twitter peaked around the time of lock-downs and social distancing in March of 2020. While the sentiment of the Tweets was initially positive, it steadily and significantly declined by late March of 2020.

### v3

@@ Abstract.
This study examines how public discourse around COVID-19 unfolded on Twitter through the lens of crisis communication and digital publics. Analyzing over 275,000 tweets involving the CDC, we identify 16 distinct discourse clusters shaped by framing, sentiment, credibility, and network dynamics. We find that CDC messaging became a flashpoint for affective and ideological polarization, with users aligning along competing frames of science vs. freedom, and public health vs. political overreach. Most clusters formed echo chambers, while a few enabled cross-cutting dialogue. Publics emerged not only around ideology but also around topical and emotional stakes, reflecting shifting concerns across different stages of the pandemic. While marginalized communities raised consistent equity concerns, these narratives struggled to reshape broader discourse. Our findings highlight the importance of long-term, adaptive engagement with diverse publics and propose design interventions—such as multi-agent AI assistants—to support more inclusive communication throughout extended public health crises.
crisis informatics, public health communication, digital publics, affective polarization, discourse analysis, temporal analysis, misinformation, marginalized communitie
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; ; $c$s: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
Twitter is a virtual “public square” (Jenzen et al., 2021) where different ad hoc “publics” can be identified using hashtags (Bruns and Moe, 2014) which are consumed by interested audiences (Livingstone, 2005). They can also be formed around mentions that are searchable by other users (Kountouri and Kollias, 2023). Retweets, quote-tweets, and responses to Tweets are associated with propagating messages in different ways.
@@
RQ1: What are the major frames of discussion in different publics and how are they polarized?
@@
RQ2: Who are the main stakeholders in our public? How are they polarized around different issues?
These disparities are not only shaped by structural inequality but also shift over time as communities move through different stages of a crisis. The evolving nature of pandemics—marked by changing levels of risk, policy, and public discourse—means that the needs of marginalized communities are not fixed. For example, initial concerns may center on access to protective equipment or credible information, while later stages may emphasize recovery support, mental health resources, or rebuilding trust in institutions. As Fox argues, the “rolling uncertainty” of emerging infectious diseases disrupts any stable ground for action or expectation (Fox, 2012), while Dalrymple cautions that mismatches between institutional narratives and lived realities can compound distrust over time (Dalrymple et al., 2016). Recognizing how marginalized communities’ concerns develop temporally is essential for shaping more responsive and equitable crisis engagement.
RQ3: How do the needs, concerns, and responses of marginalized communities evolve over the course of a pandemic, and what implications does this have for designing inclusive and adaptive public health communication strategies?
@@
We adapt the public policy discourse framing analytical framework described in Park and Lee (Park and Lee, 2020)[P.641] to model how Twitter is used as a communication channel to and from the CDC throughout a two-year public health emergency - COVID (see Figure 1). The framework suggests that public policy comprises communication between the state and society (public & private). In this framework, “communicative interaction among political actors operates in a multilayered policy landscape…[to advocate] for ”policy goals (in our case, COVID mitigation policy) and formulating the policy tools (e.g., mask and vaccine mandates) needed to achieve these goals.
We incorporate into the model three new components: (1) the affordances of propagation in digital platforms (e.g., retweets, mentions etc.); (2) message and user properties including the sentimental valence of the message, the credibility of sources, and whether the user is verified. We used these features to cluster our users into clusters representing different publics. Finally, we present a new third component representing emerging technologies, specifically, the CDC AI assistant for diverse publics presented in the discussion section (see §6.2) that can assist crisis organizations like the CDC in maintaining a line of communication with different publics. Additionally, we show how while most publics acted as echo chambers, one marginal issue public (Cluster 1) connected with the CDC and other publics. However, the model also displays the limits of this engagement by Cluster 1 as while the members of the cluster were focused on the needs of their members (two later changepoints), they could not change the trend of declining interest in racial health disparities throughout the Twitter publics (see §6.1.3).
@@ 2. Related Work
In this section, we review prior research that informs our study. We begin with work on how digital publics form and operate during crises. Next, we examine how framing and polarization influence discourse, especially in the context of public health emergencies. Finally, we consider studies on digital communication strategies in crisis response, focusing on trust, misinformation, and public engagement.
@@ 2.1. Digital Publics and Crisis Management
In these public spheres, stakeholders come together to talk about issues (c.f., (Dewey, 1927; Rauchfleisch et al., 2021)) defined as “socially constructed matters of public interest around which there is contention” (Stoltenberg, 2024, P.2009). These publics evolve over time around specific topics of interest signified by hashtags that convey the content of the messages (Bruns and Burgess, 2011).
@@ 2.1.1. Hashtags, mentions, and publics
2.2. Framing in Crisis Publics
@@
2.2.1. Misinformation and Epistemic (Dis)trust
Epistemic trust refers to the confidence placed in knowledge produced by scientists—rooted both in their technical expertise and their perceived honesty and impartiality as conveyors of scientific information (Hendriks et al., 2016; Wilholt, 2013). However, this trust has been eroding in recent years, a trend that was further exacerbated by the urgency for real-time scientific guidance during the COVID-19 pandemic (Giorgi and Eslen-Ziya, 2025). Erosion of trust in formal sources of crisis-related information has fueled the proliferation of conspiracy theories. Vulnerable groups, in particular, often exhibit heightened skepticism toward public health institutions and healthcare systems (Abelson et al., 2009). For instance, HIV/AIDS-related conspiracy theories propagated the belief that the epidemic was deliberately orchestrated by the government to harm specific populations (Kou et al., 2017). Such narratives can have serious ramifications; in this case, mistrust led some AIDS patients to avoid critical testing and treatment services (Casley, 2010). Similarly, during the Zika virus outbreak, social media emerged not only as a platform for up-to-date information but also as a breeding ground for conspiracy theories surrounding the virus (Kou et al., 2017). Kou et al. argue that these theories represented an effort by the public to interpret and cope with the disruptive societal changes brought on by the outbreak (Kou et al., 2017). This sense-making process can quickly spiral into misinformation when health authorities themselves are grappling with incomplete knowledge about the emerging threat (Gui et al., 2017). These dynamics were again evident in the discourse surrounding COVID-19 mitigation efforts (Liao et al., 2023; Vishwamitra et al., 2024; Javed et al., 2023; Wang et al., 2023).
@@ 2.3. Digital Communication in Public Health Crises
Gui et al. advocate for involving the public in crisis communication decision-making to reduce distrust between organizations like the CDC and the communities they serve (Gui et al., 2018). A notable example of this participatory approach is a COVID-19 hackathon in Germany, where participants co-developed solutions across various social domains (Haesler et al., 2021). However, public health agencies more commonly use social media as a one-way broadcast tool rather than a space for reciprocal engagement (Graham, 2014). Reframing social media as a two-way communication channel—by actively incorporating public sentiment and feedback—can help identify information gaps and overlooked concerns early, fostering greater transparency and trust.
The Crisis and Emergency Risk Communication (CERC) model emphasizes that effectively managing uncertainty during a crisis requires two-way communication between organizations and the public (Reynolds and Seeger, 2005; Renn, 2015). This approach allows the public to share feedback on crisis response efforts, fostering more responsive and transparent communication. Public-facing agencies such as the CDC and FEMA, in particular, are encouraged to adopt what Dozier et al. (Waters and Jamal, 2011) refer to as symmetrical two-way communication—a dialogic exchange that promotes mutual understanding and enables organizations to adapt their messaging strategies accordingly (Grunig and Grunig, 2008; Waters and Jamal, 2011). Despite these recommendations, prior research shows that CDC’s online communication has largely remained top-down, offering limited opportunities for public input or engagement (Gesser-Edelsburg et al., 2014). Additionally, the CERC model does not provide adequate guidance for any public health crisis that might continue for months — much like COVID (Miller et al., 2021, P.5).
This challenge is amplified in the context of emerging infectious diseases (EIDs) like COVID-19, where medical expertise and adherence to established protocols can project an “illusion of certainty” (Dalrymple et al., 2016). For instance, when CDC protocols failed to prevent the spread of Ebola, the agency responded by reinforcing existing knowledge about the disease rather than addressing the evolving nature of the threat (Dalrymple et al., 2016, P.455). Research shows that projecting certainty amid chaotic and uncertain conditions can backfire, intensifying public doubt and eroding trust in institutions like the CDC (Dalrymple et al., 2016). Pivotal events—such as the first case of disease transmission to a healthcare worker—serve as “flashpoints of change” that necessitate immediate and transparent shifts in communication strategy (Vanderford et al., 2007; Dalrymple et al., 2016).
To address these challenges, the Roadmap for Public Empowerment Policies in Crisis Management calls for adaptive and inclusive communication throughout all crisis phases—preparedness, response, and recovery. It underscores the importance of building trust prior to crises, delivering culturally sensitive and accessible messaging during crises, and actively involving communities—especially those that are marginalized or vulnerable—in recovery processes (Vos et al., 2014; Wieland et al., 2022). Key strategies include tailoring messages to specific audiences, partnering with trusted local intermediaries, and co-producing response plans with affected communities (Rowel et al., 2012).